If some eigenvalues of a matrix are repeated, the matrix is not diagonisable.true or false with full explanation

1
Expert's answer
2021-12-07T09:57:40-0500

Counterexample

Let "A=\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 2 & 0 \\
 0 & 0 & 2 \\
\end{bmatrix}." The matrix "A" is diagonal, hence diagonisable.

Find the eigenvalues of "A"


"A-\lambda I=\begin{bmatrix}
 1-\lambda & 0 & 0 \\
 0 & 2-\lambda & 0 \\
 0 & 0 & 2-\lambda \\
\end{bmatrix}"

"\det(A-\lambda I)=\begin{vmatrix}
 1-\lambda & 0 & 0 \\
 0 & 2-\lambda & 0 \\
 0 & 0 & 2-\lambda \\
\end{vmatrix}"

"=(1-\lambda)(2-\lambda)^2=0"

"\lambda_1=1, \lambda_2=\lambda_3=2."

The matrix "A" has eigenvalues "1,2,2" (not all distinct), but is diagonisable.


False.